Serpens South Trial DSL11Flags
2012-11-03 c0924I.4D_88Serpen.SL.11 5.5 A- Comments: NODE_NOT_FOUND [2012-11-03 22:54]jmc : Scripted was restarted due to bug in obsdef3 script (which will not otherwise affect the data). Ant 6 offline because telescope drive controller is not working. 2012-11-03 c0924I.4D_88Serpen.WB.11 5.5 C+ Comments: NODE_NOT_FOUND [2012-11-03 22:57]jmc : Gave track a C+ grade to make sure time is not counted. data look fine. C6 offline because of had drive controller box.
* Emailed Nikolaus about C18 and C19 gains consistently ~1.3 * cont rms: box1~7.4 mJy/bm ; box2~8 mJy/bm * HCO+ single channel rms: box1~564 mJy/bm ; box2~650 mJy/bm
* c6 out. * C16 also has higher tsys. C18 & C19 are the outriggers... is that a coincidence? * RMS cont, box1/2: 6.8/7 mJy/b * RMS HCO+, box1/2: 1.0/1.1 Jy/b * SD: seems OK although weak signal * WB: -bline(1,4): amplitude ripple -1743-038 flux: measured=4.08, adopted= 3.6 (why this difference?) -rms: 2 mJy/b
select='ant(2),time(21:31:25.5,22:09:52.5)' #high gain select='ant(12),time(23:33:22.5,00:22:07)' #high gain select='ant(16),time(00:39:59,00:43:14)' #tsys
select="ant(2),time(21:53,21:56)" # high gain select="ant(12),time(23:55,23:59)" # high gain select="ant(16),time(02:27,02:31)" # high gain